Four point charges arw equally separated by 1.00m in air as shown.If q1=3.0nC,q2=4.0nC,q3=2.0nC,q4=5.0nC,what is the electric field at a point halfway between q2 and q3?


1
Expert's answer
2022-02-04T07:41:59-0500

The electric field at a point halfway between q2 and q3 is


"E=\dfrac{1}{4\pi\epsilon_0}\bigg(\dfrac{q_1}{1.5^2}+\dfrac{q_2}{0.5^2}-\dfrac{q_3}{0.5^2}-\dfrac{q_4}{1.5^2}\bigg),\\\space\\
E=63.9\text{ V/m}."